# Structured Sampling for Robust Euclidean Distance Geometry

### v1

@@ I Introduction
In certain scenarios, missing distances exhibit a structured pattern. For instance, in the sensor localization problem [20, 21], there are typically two types of nodes: anchor nodes with known locations and target nodes with unknown locations. A standard setup involves knowing the distances between the target nodes and all anchors while lacking information about distances between the target nodes themselves. Motivated by such setups, a recent work in [22] has developed a theoretical framework for recovering the locations of sensor nodes when provided with partial exact distance information between anchors and partial exact distance information between target nodes and anchors. In this paper, we consider a sampling model inspired by the standard sensor localization problem. Specifically, we address the scenario where distances between target nodes and anchors are corrupted by outliers, while distances between anchors are assumed to be exact. We propose a novel algorithm for this problem, powered by Nyström method and robust principal component analysis (PCA).
@@ I-B Notation
In this paper, we use the following notations. Uppercase boldface scripts, such as $\bm{M}$, denote matrices, while lowercase boldface scripts, such as $\bm{v}$, represent vectors. The $(i,j)$-th entry of a matrix $\bm{M}$ is denoted by $[\bm{M}]_{i,j}$. The transpose of a matrix $\bm{M}$ is denoted by $\bm{M}^{\top}$, and the pseudo-inverse of $\bm{M}$ is represented as $\bm{M}^{\dagger}$. The standard $\ell_{2}$-norm is denoted by $\|\cdot\|_{2}$. The vector of ones of size $m$ is denoted by $\bm{1}_{m}$, and the matrix of ones of size $m\times n$ is denoted by $\bm{1}_{m\times n}$. Finally, $\mu_{(\bm{M})}$ denotes the mean of all elements of the matrix $\bm{M}$.
@@ II-B Robust PCA
Robust Principal Component Analysis (RPCA) is a fundamental machine learning tool for recovering a low-rank data matrix from sparse outlier corruptions. That is, given a corrupted observation $\bm{Y}=\bm{L}+\bm{S}$ where $\bm{L}$ is low-rank and $\bm{S}$ is sparse, RPCA recovers $\bm{L}$ and $\bm{S}$ simultaneously. RPCA is widely used in applications [37, 38, 39, 40, 41, 42, 43, 44, 45], and numerous convex and non-convex solvers have been extensively studied under various settings [46, 47, 48, 49, 50, 51, 52, 53, 54].
In this work, we employ RPCA to eliminate outliers from the $\bm{F}$ block of the distance matrix $\bm{D}$. Among various existing RPCA solvers, we use the non-convex accelerated alternating projections (AccAltProj) method [49], which is robust, lightweight, and highly efficient: It has a recovery guarantee with linear convergence. Its main drawback is requiring the knowledge of the exact rank of $\bm{L}$; however, this is not an issue for EDG problems since the true rank is explicit as discussed in Section I-A.
@@ III-A Nyström approximation for distance matrices
where $\bm{s}$ is a vector with entries that sum to 1. This operation generates a family of valid solutions for $\bm{X}$ based on different choices of $s$. Following [57], we select $s$ with entries equal to $1/m$ for the first $m$ positions and 0 elsewhere. In this paper, we propose a robust EDG framework that leverages non-orthogonal measurements to estimate the Gram matrix based on the relationship defined in (3). The blocks $\bm{A}$ and $\bm{B}$ in (3) are computed as follows:
$\bm{A}$ $=-\frac{1}{2}\left(\bm{E}-\frac{1}{m}\bm{E}\bm{1}_{m\times m}-\frac{1}{m}\bm{1}_{m\times m}\bm{E}+\mu_{(\bm{E})}\bm{1}_{m\times m}\right),$ (4)
$\bm{B}$ $=-\frac{1}{2}\left(\bm{F}-\frac{1}{m}\bm{1}_{m\times m}\bm{F}-\frac{1}{m}\bm{E}\bm{1}_{m\times n}+\mu_{(\bm{E})}\bm{1}_{m\times n}\right).$ (5)
@@
which is exact when $\operatorname{rank}(\bm{A})=\operatorname{rank}(\bm{X})$ [58].
@@ IV Proposed approach
Once the sparse noise is removed, we estimate the block $\bm{B}$ of the Gram matrix $\bm{X}$ from the clean $\hat{\bm{F}}$ and the available $\bm{E}$ block of $\bm{D}$. The reconstruction equations (4) and (5) are derived from the relationship between squared distance matrices and Gram matrices. The estimation of $\bm{B}$ relies only on the $\bm{E}$ and $\hat{\bm{F}}$ blocks, making the process efficient and independent of the missing $\bm{G}$ block. The Nyström approximation is then applied to estimate the missing $\bm{C}$ block of the Gram matrix $\bm{X}$, as described in (6). Finally, the reconstructed blocks $\hat{\bm{B}}$, and $\bm{C}$ along with $\bm{A}$ are assembled to obtain the estimate of the Gram matrix $\bm{X}$. The complete reconstruction algorithm, Structured Robust EDG (SREDG), is summarized in Algorithm 1.
Algorithm 1 Structured Robust EDG (SREDG)
1:Input: $\bm{E}\in{\mathbb{R}}^{m\times m}$: squared distance matrix among anchor nodes; $\bm{F}\in{\mathbb{R}}^{n\times n}$: noisy squared distance matrix among anchor and target nodes; $r$: rank of $\bm{D}$; $\operatorname{RPCA}$: the chosen RPCA solver.
@@
6:$\hat{\bm{X}}=\begin{bmatrix}\bm{A}&\hat{\bm{B}}\\ \hat{\bm{B}}^{\top}&\hat{\bm{C}}.\end{bmatrix}$
@@
We project the estimate $\hat{\bm{X}}$ onto the set of rank-$r$ positive semi-definite matrices via truncated eigenvalue decomposition ${\hat{\bm{U}}}_{r}\hat{\bm{\Lambda}}_{r}\hat{\bm{U}}_{r}^{\top}$ and set any negative eigenvalues to 0. The estimated coordinates are then computed as $\hat{\bm{P}}=\hat{\bm{\Lambda}}_{r}^{\frac{1}{2}}\hat{\bm{U}}_{r}^{\top}$. This step yields the $r$-dimensional Euclidean embedding of the nodes centered at the origin. Note that the embedding is unique up to rigid motions.
@@ V-A Synthetic data experiments
To recover the true distances from the corrupted measurements, we employ SREDG (i.e., Algorithm 1). A truncated eigendecomposition followed by Procrustes analysis [63] is used to compute the root mean square error ($\mathrm{RMSE}$), which is defined as:
$\mathrm{RMSE}^{2}={\frac{1}{T}\sum_{i=1}^{T}\|\hat{\bm{p}}_{i}-\bm{p}_{i}\|_{2}^{2}}.$
@@ VI Conclusion
In conclusion, this paper introduces a novel algorithm, Structured Robust EDG (SREDG), designed for estimating the positions of points with sparse outliers in the distance measurements between anchor and target nodes. The algorithm’s effectiveness is demonstrated through empirical evaluations on synthetic sensor localization datasets and molecular experiments, showing that it performs well even with a limited number of anchor nodes and in the presence of substantial sparse noise. Future work will focus on establishing provable error bounds for point reconstruction in SREDG.

### v2

@@ I Introduction
In certain scenarios, missing distances exhibit a structured pattern. For instance, in the sensor localization problem [20, 21], there are typically two types of nodes: anchor nodes with known locations and target nodes with unknown locations. A standard setup involves knowing the distances between the target nodes and all anchors while lacking information about distances between the target nodes themselves. Motivated by such setups, a recent work in [22] has developed a theoretical framework for recovering the locations of sensor nodes when provided with partial exact distance information between anchors and partial exact distance information between target nodes and anchors. In this paper, we consider a sampling model inspired by the standard sensor localization problem. Specifically, we address the scenario where distances between target nodes and anchors are corrupted by outliers, while distances between anchors are assumed to be exact. We propose a novel algorithm for this problem, powered by Nyström method and robust principal component analysis (RPCA).
@@ I-B Notation
In this paper, we use the following notations. Uppercase boldface scripts, such as $\bm{M}$, denote matrices, while lowercase boldface scripts, such as $\bm{v}$, represent vectors. The $(i,j)$-th entry of a matrix $\bm{M}$ is denoted by $[\bm{M}]_{i,j}$. The transpose of a matrix $\bm{M}$ is denoted by $\bm{M}^{\top}$, and the pseudo-inverse of $\bm{M}$ is represented as $\bm{M}^{\dagger}$. The standard $\ell_{2}$-norm is denoted by $\|\cdot\|_{2}$. The vector of ones of size $m$ is denoted by $\bm{1}_{m}$, and the matrix of ones of size $m\times n$ is denoted by $\bm{1}_{m\times n}$. Finally, $\mathbb{E}([\bm{M}]_{i,j})$ denotes the mean of all elements of the matrix $\bm{M}$.
@@ II-B Robust PCA
Robust Principal Component Analysis (RPCA) is a fundamental machine learning tool for recovering a low-rank data matrix from sparse outlier corruptions. That is, given a corrupted observation $\bm{Y}=\bm{L}+\bm{S}$ where $\bm{L}$ is low-rank and $\bm{S}$ is sparse, RPCA recovers $\bm{L}$ and $\bm{S}$ simultaneously. RPCA is widely used in applications [37, 38, 39, 40, 41, 42, 43, 44, 45], and numerous convex and non-convex solvers have been extensively studied under various settings [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
In this work, we employ RPCA to eliminate outliers from the $\bm{F}$ block of the distance matrix $\bm{D}$. Among various existing RPCA solvers, we use the non-convex Accelerated Alternating Projections (AccAltProj) algorithm [49], which is robust, lightweight, and highly efficient. AccAltProj has a recovery guarantee with linear convergence. Its main drawback is requiring the knowledge of the exact rank of $\bm{L}$; however, this is not an issue for EDG problems since the true rank is explicit as discussed in Section I-A.
@@ III-A Nyström approximation for distance matrices
where $\bm{s}$ is a vector with entries that sum to 1. This operation generates a family of valid solutions for $\bm{X}$ based on different choices of $\bm{s}$. Following [59], we select $\bm{s}$ with entries equal to $1/m$ for the first $m$ positions and 0 elsewhere. In this paper, we propose a robust EDG framework that leverages non-orthogonal measurements to estimate the Gram matrix based on the relationship defined in (3). The blocks $\bm{A}$ and $\bm{B}$ in (3) are computed as follows:
$\bm{A}$ $=-\frac{1}{2}\left(\bm{E}-\frac{1}{m}\bm{E}\bm{1}_{m\times m}-\frac{1}{m}\bm{1}_{m\times m}\bm{E}+\mathbb{E}([\bm{M}]_{i,j})\bm{1}_{m\times m}\right),$ (4)
$\bm{B}$ $=-\frac{1}{2}\left(\bm{F}-\frac{1}{m}\bm{1}_{m\times m}\bm{F}-\frac{1}{m}\bm{E}\bm{1}_{m\times n}+\mathbb{E}([\bm{M}]_{i,j})\bm{1}_{m\times n}\right).$ (5)
@@
which becomes exact when $\operatorname{rank}(\bm{A})=\operatorname{rank}(\bm{X})$ [60].
@@ IV Proposed approach
Once the sparse noise is removed, we estimate the block $\bm{B}$ of the Gram matrix $\bm{X}$ from the clean $\hat{\bm{F}}$ and the available $\bm{E}$ block of $\bm{D}$. The reconstruction equations (4) and (5) are derived from the relationship between squared distance matrices and Gram matrices. The estimation of $\bm{B}$ relies only on the $\bm{E}$ and $\hat{\bm{F}}$ blocks, making the process efficient and independent of the missing $\bm{G}$ block. The Nyström approximation is then applied to estimate the missing $\bm{C}$ block of the Gram matrix $\bm{X}$, as described in (6). Finally, the reconstructed blocks $\hat{\bm{B}}$, and $\bm{C}$ along with $\bm{A}$ are assembled to obtain the estimate of the Gram matrix $\bm{X}$. The complete reconstruction algorithm, namely Structured Robust EDG, is summarized in Algorithm 1.
Algorithm 1 Structured Robust EDG
1:Input: $\bm{E}\in{\mathbb{R}}^{m\times m}$: squared distance matrix among anchor nodes; $\bm{F}\in{\mathbb{R}}^{n\times n}$: noisy squared distance matrix among anchor and target nodes; $\bar{r}=r+2$: rank of $\bm{D}$; $\operatorname{RPCA}$: the chosen RPCA solver.
@@
6:$\hat{\bm{X}}=\begin{bmatrix}\bm{A}&\hat{\bm{B}}\\ \hat{\bm{B}}^{\top}&\hat{\bm{C}}\end{bmatrix}$.
@@
We project the estimate $\hat{\bm{X}}$ onto the set of rank-$r$ positive semi-definite matrices via truncated eigenvalue decomposition ${\hat{\bm{U}}}_{r}\hat{\bm{\Lambda}}_{r}\hat{\bm{U}}_{r}^{\top}$, where $r=\bar{r}-2$, and set any negative eigenvalues to 0. The estimated coordinates are then computed as $\hat{\bm{P}}=\hat{\bm{\Lambda}}_{r}^{\frac{1}{2}}\hat{\bm{U}}_{r}^{\top}$. This step yields the $r$-dimensional Euclidean embedding of the nodes centered at the origin. Note that the embedding is unique up to rigid motions.
@@ V-A Synthetic data experiments
To recover the true distances from the corrupted measurements, we employ Structured Robust EDG (i.e., Algorithm 1). A truncated eigendecomposition followed by Procrustes analysis [65] is used to compute the root mean square error ($\mathrm{RMSE}$), which is defined as:
$\mathrm{RMSE}^{2}={\frac{1}{T}\sum_{i=1}^{T}\left\|\hat{\bm{p}}_{i}-\bm{p}_{i}\right\|_{2}^{2}}.$
@@ VI Conclusion
In conclusion, this paper introduces a novel algorithm, dubbed Structured Robust EDG, designed for estimating the positions of points with sparse outliers in the distance measurements between anchor and target nodes. The algorithm’s effectiveness is demonstrated through empirical evaluations on synthetic sensor localization datasets and molecular experiments, showing that it performs well even with a limited number of anchor nodes and in the presence of substantial sparse noise. Future work will focus on establishing provable error bounds for point reconstruction in Structured Robust EDG.